### v1

@@ Abstract
We ask whether a stationary lattice in dimension $d$ whose points are shifted by identically distributed but possibly dependent perturbations remains hyperuniform. When $d=1$ or $2$, we show that it is the case when the perturbations have a finite $d$-moment, and that this condition is sharp. When $d\geq 3$, we construct arbitrarily small perturbations such that the resulting point process is not hyperuniform. As a side remark of independent interest, we exhibit hyperuniform processes with arbitrarily slow decay of their number variance.
@@ 1.1 Setting and main results
The basic example of a class I-hyperuniform point process is the “stationary lattice” $\mathbf{L}:=\{x+U,\ x\in\mathcal{L}\}$, where $\mathcal{L}$ is some lattice in $\mathbb{R}^{\mathsf{d}}$ (i.e. $\mathbb{Z}^{\mathsf{d}}$ or the image of $\mathbb{Z}^{\mathsf{d}}$ by an affine transformation), and the random variable $U$ is uniformly distributed in a fundamental domain of $\mathcal{L}$. Since (1.2) is the slowest possible growth, stationary lattices are “as hyperuniform as it gets”. In this paper, we consider perturbed lattices of the form $\mathbf{PL}:=\{x+U+\mathrm{p}_{x},\ x\in\mathcal{L}\}$ where $U$ is as above while $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ is a family of random variables independent of $U$, thought of as “perturbations”, with values in $\mathbb{R}^{\mathsf{d}}$, and we discuss whether $\mathbf{PL}$ remains hyperuniform or not.
When the perturbations are independent and identically distributed, the answer is well-known: in any dimension, if $\mathrm{p}_{0}$ has a finite first moment, then $\mathbf{PL}$ remains hyperuniform of class I, see [GS75]. In fact, even if $\mathrm{p}_{0}$ has only a finite moment of order $10<s<$ 1, then $\mathbf{PL}$ remains hyperuniform (but not necessarily of class I).
@@
We thus obtain with the first item a sharp sufficient condition for class-I hyperuniformity of perturbed lattices in dimension $1$ based only on the size of the perturbations. The second item indicates that in dimension $2$ (or higher) this is not possible - any sufficient condition should presumably involve the correlation structure of the perturbations, too (see Section 1.5).
@@ 1.4 An optimal transport perspective
where $\mathsf{cpl}(\xi,\eta)$ denotes the set of all couplings between the point configurations $\xi$ and $\eta$, seen as measures on $\mathbb{R}^{\mathsf{d}}$. The quantity $c_{p}(\xi,\eta)$ can be thought of as a “$p$-Wasserstein distance per unit volume” between the measures $\xi$ and $\eta$.
@@
$\mathrm{Wass}_{p}^{p}(\mathsf{P}_{0},\mathsf{P}_{1})=\inf_{(\xi^{\bullet},\eta^{\bullet})}\inf_{q^{\bullet}\in\mathsf{cpl}_{e}(\xi^{\bullet},\eta^{\bullet})}\mathbb{E}\left[\int_{[0,1]^{\mathsf{d}}\times\mathbb{R}^{d}}|x-y|^{p}\mathsf{d}q(x.y)\right],$ (1.5)
@@
With this language, the results of Theorem 1 and Theorem 2 can be rephrased as follows:
@@ 1.6 Plan of the paper
The analogous positive result in dimension $\mathsf{d}=1$ is simpler to prove and is provided in Appendix B, along with the proof of the second item of Theorem 2.
@@ 2.4 Basic properties of perturbed lattices
It is straightforward to check that, in the stationary case, those two notions coincide.
@@ 3.1 Preliminaries
    $\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\right)\right|\leq g(t)\min\left(|t|^{2},1\right),\text{ for all }t\in\mathbb{R}^{2}.$ (3.2)
@@
$\left|\mathbb{E}[e^{-is\cdot X}]-\sum_{k=0}^{n}\frac{(-is)^{k}}{k!}\mathbb{E}[X^{k}]\right|\leq\mathbb{E}\left[\min(|sX|^{n+1},2|sX|^{n})\right]\text{ for all $s\in\mathbb{R}^{\mathsf{d}}$.}$ (3.3)
@@
$\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\right)\right|\preceq|t|^{2}\mathbb{E}\left[\min\left(|t||\mathrm{p}_{x}-\mathrm{p}_{0}|^{3},|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right)\right].$
@@
$\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)=1-\frac{|t|^{2}}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]+o(|t|^{2}),$ (3.5)
@@
Truncation near the origin.
@@
Decomposing the term $B$.
Truncation.
@@
$\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)\chi_{\ell}(t)\mathtt{d}t=-\frac{1}{2}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t\\ +\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)o(|t|^{2})\chi_{\ell}(t)\mathtt{d}t.$
@@
$B_{1}:=-\frac{1}{2}\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t,$ (3.10)
@@ 3.2 The main term $B_{1}$
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L},|x|\leq 100r}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t.$ (3.13)
@@
Re-introducing some terms in the sum.
@@
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L}}\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)|t|^{2}\chi_{\ell}(t)\mathtt{d}t+o_{r}(1),$ (3.15)
with an error term bounded by $\mathcal{O}((\ell r)^{-1})$.
The spectral measure.
Let $g_{\ell}$ be the following map, appearing in the integrand in (3.13):
$g_{\ell}:t\mapsto\chi_{\ell}(t)|t|^{2}.$
Introduce the autocorrelation function of the perturbation field $(\mathrm{p}_{x})_{x\in\mathcal{L}}$, namely the function $G:\mathcal{L}\to\mathbb{R}$ defined by (recall that the perturbations are assumed to be centered):
$G(x):=\mathbb{E}[\mathrm{p}_{x}\cdot\mathrm{p}_{0}].$ (3.16)
@@
  • The autocorrelation function is bounded by $|G(x)|\leq G(0)$ for all $x\in\mathcal{L}$.
  • $G$ is a positive-definite function on $\mathcal{L}$. In particular, by Bochner’s theorem (see [Fol16, Thm. (4.18)]), there exists a “spectral measure” $\rho$ which is a non-negative finite measure on the Pontryagin dual $\Omega$ of the locally compact abelian group $\mathcal{L}$, such that:
    $G(x)=\int_{\Omega}e^{-2i\pi x\cdot\omega}\mathtt{d}\rho(\omega)\text{ for all }x\in\mathcal{L}.$ (3.17)
@@
Since $\mathbb{E}\left[|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right]=2(G(0)-G(x))$ for all $x\in\mathcal{L}$, we may return to (3.15), use our notation, and write $B_{1}$ as $B_{1}=B_{11}-B_{12}+o_{r}(1)$, with:
$B_{11}:=\sum_{x\in\mathcal{L}}G(x)\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t,\quad B_{12}:=G(0)\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t.$
Controlling $B_{11}.$
Using the spectral measure $\rho$ as in (3.17), we can re-write $B_{11}$ as:
$B_{11}=\sum_{x\in\mathcal{L}}\int_{\Omega}e^{-2i\pi x\cdot\omega}\left(\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t\right)\ \mathtt{d}\rho(\omega)=\int_{\Omega}\sum_{x\in\mathcal{L}}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}}(x)\mathtt{d}\rho(\omega).$
Applying Poisson’s formula (2.1) to the sum $\sum_{x\in\mathcal{L}}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}}(x)$ for all fixed $\omega\in\Omega$, we obtain:
$B_{11}=2\pi\sum_{k\in\mathcal{L}^{\star}}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}(2\pi k+2\pi\omega)\mathtt{d}\rho(\omega).$ (3.18)
Note that $g_{\ell}$ is a Schwartz function and thus so is $\widehat{g_{\ell}}$ and (since $\widehat{j_{r}}$ is $L^{1}$), so is $\widehat{j_{r}\times g_{\ell}}$ , which justifies our interversions of sums over $x\in\mathcal{L}$ and integrals against $\mathtt{d}\rho$, as well as the use of Poisson’s formula. We now distinguish between the lattice points $k=0$ and $k\in\mathcal{L}^{\star}\setminus\{0\}$.
For $k\neq 0$.
@@
$B_{12}:=G(0)\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)|t|^{2}\mathtt{d}t=G(0)\sum_{x\in\mathcal{L}}\widehat{j_{r}\times g_{l}}(x)=2\pi G(0)\sum_{k\in\mathcal{L}^{\star}}j_{r}(2\pi k)g_{l}(2\pi k)=0.$
Indeed $j_{r}(0)g_{l}(0)=0$, and since $g_{\ell}$ is supported on $\mathrm{B}_{\ell}$ the other summands vanish for $\ell$ small enough (depending on $\mathcal{L}$) thus for $r$ large enough. ∎
Conclusion.

### v2

@@ Abstract
We ask whether a stationary lattice in dimension $\mathsf{d}$ whose points are shifted by identically distributed but possibly dependent perturbations remains hyperuniform. When $\mathsf{d}=1$ or $2$, we show that it is the case when the perturbations have a finite $\mathsf{d}$-moment, and that this condition is sharp. When $\mathsf{d}\geq 3$, we construct arbitrarily small perturbations such that the resulting point process is not hyperuniform. As a side remark of independent interest, we exhibit hyperuniform processes with arbitrarily slow decay of their number variance.
@@ 1.1 Setting and main results
The basic example of a class I-hyperuniform point process is the “stationary lattice” $\mathbf{L}:=\{x+U,\ x\in\mathcal{L}\}$, where $\mathcal{L}$ is some lattice in $\mathbb{R}^{\mathsf{d}}$ (i.e. $\mathbb{Z}^{\mathsf{d}}$ or the image of $\mathbb{Z}^{\mathsf{d}}$ by an affine transformation), and the random variable $U$ is uniformly distributed in a fundamental domain of $\mathcal{L}$. Since (1.2) is the slowest possible growth (cf. [Cos21, Thm. 4.2], [Bec87, Thm. 2A]), stationary lattices are “as hyperuniform as it gets”. In this paper, we consider perturbed lattices of the form $\mathbf{PL}:=\{x+U+\mathrm{p}_{x},\ x\in\mathcal{L}\}$ where $U$ is as above while $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ is a family of random variables independent of $U$, thought of as “perturbations”, with values in $\mathbb{R}^{\mathsf{d}}$, and we discuss whether $\mathbf{PL}$ remains hyperuniform or not.
When the perturbations are independent and identically distributed, the answer is well-known: in any dimension, if $\mathrm{p}_{0}$ has a finite first moment, then $\mathbf{PL}$ remains hyperuniform of class I, see [GS75]. In fact, even if $\mathrm{p}_{0}$ has only a finite moment of order $10<s<$ 1, then $\mathbf{PL}$ remains hyperuniform (but not necessarily of class I). In the setting of dependent field $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ with (or without) information about the structure of dependence, the question was mainly open whereas many models are involved by this setting [Fli25, Gab04, HL24, KLLY25].
@@
We thus obtain with the first item a sharp sufficient condition for class-I hyperuniformity of perturbed lattices in dimension $1$ based only on the size of the perturbations. The sharpness means that we can construct a perturbation such that $\mathbb{E}[|\mathrm{p}_{0}|^{2}]=+\infty$ and $\mathbf{PL}$ is not class I hyperuniform, see Remark 4.3. The second item indicates that in dimension $2$ (or higher) this is not possible - any sufficient condition should presumably involve the correlation structure of the perturbations, too (see Section 1.5).
@@ 1.4 An optimal transport perspective
where $\mathsf{cpl}(\xi,\eta)$ denotes the set of all couplings between the point configurations $\xi$ and $\eta$, seen as measures on $\mathbb{R}^{\mathsf{d}}$ and $\Lambda_{n}$ is the box $[-n/2,n/2]^{d}$. The quantity $c_{p}(\xi,\eta)$ can be thought of as a “$p$-Wasserstein distance per unit volume” between the measures $\xi$ and $\eta$.
@@
$\mathrm{Wass}_{p}^{p}(\mathsf{P}_{0},\mathsf{P}_{1})=\inf_{(\xi^{\bullet},\eta^{\bullet})}\inf_{q^{\bullet}\in\mathsf{cpl}_{e}(\xi^{\bullet},\eta^{\bullet})}\mathbb{E}\left[\int_{[0,1]^{\mathsf{d}}\times\mathbb{R}^{d}}|x-y|^{p}q(dx,dy)\right],$ (1.5)
@@
With this language, the main results of Theorem 1 and Theorem 2 can be rephrased as follows:
@@ 1.6 Plan of the paper
The analogous positive result in dimension $\mathsf{d}=1$ is simpler to prove and is provided in Appendix B, along with the proof of the first item of Theorem 2.
@@ 2.4 Basic properties of perturbed lattices
It is straightforward to check that, when the perturbations themselves are stationary, those two notions coincide.
@@ 3.1 Preliminaries
    $\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{1}{2}\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\right)\right|\leq g(t)\min\left(|t|^{2},1\right),\text{ for all }t\in\mathbb{R}^{2}.$ (3.2)
@@
$\left|\mathbb{E}[e^{-is\cdot X}]-\sum_{k=0}^{n}\frac{(-i)^{k}}{k!}\mathbb{E}[(s\cdot X)^{k}]\right|\leq\mathbb{E}\left[\min(|s\cdot X|^{n+1},2|s\cdot X|^{n})\right]\text{ for all $s\in\mathbb{R}^{\mathsf{d}}$.}$ (3.3)
@@
$\left|\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-\left(1-\frac{1}{2}\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\right)\right|\preceq|t|^{2}\mathbb{E}\left[\min\left(|t||\mathrm{p}_{x}-\mathrm{p}_{0}|^{3},|\mathrm{p}_{x}-\mathrm{p}_{0}|^{2}\right)\right].$
@@
$\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)=1-\frac{1}{2}\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]+o(|t|^{2}),$ (3.5)
@@
Truncation near the origin
@@
Decomposing the term $B$
Truncation
@@
$\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\left(\varphi_{\mathrm{p}_{x}-\mathrm{p}_{0}}(t)-1\right)\chi_{\ell}(t)\mathtt{d}t=-\frac{1}{2}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\chi_{\ell}(t)\mathtt{d}t\\ +\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)o(|t|^{2})\chi_{\ell}(t)\mathtt{d}t.$
@@
$B_{1}:=-\frac{1}{2}\sum_{x\in\mathcal{L}\setminus\{0\},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\chi_{\ell}(t)\mathtt{d}t,$ (3.10)
@@ 3.2 The main term $B_{1}$
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L},|x|\leq 100r}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\chi_{\ell}(t)\mathtt{d}t.$ (3.13)
@@
Re-introducing some terms in the sum
@@
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\mathbb{E}\left[\big{(}t\cdot(\mathrm{p}_{x}-\mathrm{p}_{0})\big{)}^{2}\right]\chi_{\ell}(t)\mathtt{d}t+o_{r}(1),$ (3.15)
with an error term bounded by $\mathcal{O}((\ell r)^{-1})$. Let us now expand the scalar product in the expectation by introducing coordinates $t=(t_{1},t_{2})$ and $\mathrm{p}_{x}=(\mathrm{p}_{x,1},\mathrm{p}_{x,2})$:
$B_{1}=-\frac{1}{2}\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\Big{(}\mathbb{E}\left[(\mathrm{p}_{x,1}-\mathrm{p}_{0,1})^{2}\right]\chi_{\ell}(t)t_{1}^{2}\\ +\mathbb{E}\left[(\mathrm{p}_{x,2}-\mathrm{p}_{0,2})^{2}\right]\chi_{\ell}(t)t_{2}^{2}+2\mathbb{E}\left[(\mathrm{p}_{x,1}-\mathrm{p}_{0,1})(\mathrm{p}_{x,2}-\mathrm{p}_{0,2})\right]\chi_{\ell}(t)t_{1}t_{2}\Big{)}\mathtt{d}t+o_{r}(1).$ (3.16)
The spectral measure
Introduce the following maps, as in the integrand in (3.16):
$g_{\ell}^{1}:t\mapsto\chi_{\ell}(t)t_{1}^{2},\quad g_{\ell}^{2}:t\mapsto\chi_{\ell}(t)t_{2}^{2}\quad\text{and}\quad g_{\ell}^{1,2}:t\mapsto\chi_{\ell}(t)t_{1}t_{2}.$
We also introduce several variants of autocorrelation functions for the perturbation field $(\mathrm{p}_{x})_{x\in\mathcal{L}}$ (recall that the perturbations are assumed to be centered): for $x\in\mathcal{L}$ we let
$G^{1}(x):=\mathbb{E}[\mathrm{p}_{x,1}\cdot\mathrm{p}_{0,1}]\quad\text{ and}\quad G^{2}(x):=\mathbb{E}[\mathrm{p}_{x,2}\cdot\mathrm{p}_{0,2}]$
as well as
$G^{1,2+}(x):=\mathbb{E}[(\mathrm{p}_{x,1}+\mathrm{p}_{x,2})\cdot(\mathrm{p}_{0,1}+\mathrm{p}_{0,2})]\quad\text{ and}\quad G^{1,2-}(x):=\mathbb{E}[(\mathrm{p}_{x,1}-\mathrm{p}_{x,2})\cdot(\mathrm{p}_{0,1}-\mathrm{p}_{0,2})].$
@@
  • These autocorrelation functions are bounded by their values at 00.
  • Both $G^{1}$, $G^{2}$ and $G^{1,2\pm}$ are positive-definite functions on $\mathcal{L}$. In particular, by Bochner’s theorem (see [Fol16, Thm. (4.18)]), there exist “spectral measures” $\rho^{1},\rho^{2},\rho^{1,2\pm}$ which are non-negative finite measures on the Pontryagin dual $\Omega$ of the locally compact abelian group $\mathcal{L}$, such that for all $x\in\mathcal{L}$ and $m=1,2$
    $G^{m}(x)=\int_{\Omega}e^{-2i\pi x\cdot\omega}\mathtt{d}\rho^{m}(\omega)\quad\text{and}\quad G^{1,2\pm}(x)=\int_{\Omega}e^{-2i\pi x\cdot\omega}\mathtt{d}\rho^{1,2\pm}(\omega).$ (3.17)
@@
We can express the expectations in (3.16) using these autocorrelation functions by noting first that $\mathbb{E}\left[(\mathrm{p}^{m}_{x}-\mathrm{p}^{m}_{0})^{2}\right]=2(G^{m}(0)-G^{m}(x))$ for $m=1,2$ and that moreover:
$\mathbb{E}\left[(\mathrm{p}_{x,1}-\mathrm{p}_{0,1})(\mathrm{p}_{x,2}-\mathrm{p}_{0,2})\right]$ $=$ $\frac{1}{4}\mathbb{E}\left[\Big{(}(\mathrm{p}_{x,1}-\mathrm{p}_{0,1})+(\mathrm{p}_{x,2}-\mathrm{p}_{0,2})\Big{)}^{2}-\Big{(}(\mathrm{p}_{x,1}-\mathrm{p}_{0,1})-(\mathrm{p}_{x,2}-\mathrm{p}_{0,2})\Big{)}^{2}\right]$
$=$ $\frac{1}{4}\mathbb{E}\left[\Big{(}(\mathrm{p}_{x,1}+\mathrm{p}_{x,2})-(\mathrm{p}_{0,1}+\mathrm{p}_{0,2})\Big{)}^{2}-\Big{(}(\mathrm{p}_{x,1}-\mathrm{p}_{x,2})-(\mathrm{p}_{0,1}-\mathrm{p}_{0,2})\Big{)}^{2}\right]$
$=$ $\frac{1}{2}\left(G^{1,2+}(0)-G^{1,2+}(x)-G^{1,2-}(0)+G^{1,2-}(x)\right).$
We may now return to (3.16), use our notation, and write $B_{1}$ as $B_{1}=B_{11}-B_{12}+o_{r}(1)$, with:
$B_{11}:=\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)\left(G^{1}(x)t_{1}^{2}+G^{2}(x)t_{2}^{2}+\frac{1}{2}(G^{1,2+}(x)-G^{1,2-}(x))t_{1}t_{2}\right)\mathtt{d}t$
and
$B_{12}:=\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)\left(G^{1}(0)t_{1}^{2}+G^{2}(0)t_{2}^{2}+\frac{1}{2}(G^{1,2+}(0)-G^{1,2-}(0))t_{1}t_{2}\right)\mathtt{d}t.$
Controlling $B_{11}$
Using the spectral measures $\rho^{1},\rho^{2}$ and $\rho^{1,2\pm}$ as in (3.17), we can re-write $B_{11}$ as:
$B_{11}$ $=$ $\sum_{x\in\mathcal{L}}\int_{\Omega}e^{-2i\pi x\cdot\omega}\left(\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)t_{1}^{2}\mathtt{d}t\right)\mathtt{d}\rho^{1}(\omega)$
$+\int_{\Omega}e^{-2i\pi x\cdot\omega}\left(\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)t_{2}^{2}\mathtt{d}t\right)\mathtt{d}\rho^{2}(\omega)$
$+\frac{1}{2}\int_{\Omega}e^{-2i\pi x\cdot\omega}\left(\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)t_{1}t_{2}\mathtt{d}t\right)\ \mathtt{d}(\rho^{1,2+}-\rho^{1,2-})(\omega)$
$=$ $\sum_{x\in\mathcal{L}}\int_{\Omega}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}^{1}}(x)\ \mathtt{d}\rho^{1}(\omega)+\int_{\Omega}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}^{2}}(x)\ \mathtt{d}\rho^{2}(\omega)$
$+\frac{1}{2}\int_{\Omega}e^{-2i\pi x\cdot\omega}\widehat{j_{r}\times g_{\ell}^{1,2}}(x)\ \mathtt{d}(\rho^{1,2+}-\rho^{1,2-})(\omega)$
Applying Poisson’s formula (2.1) for all fixed $\omega\in\Omega$, we obtain:
$B_{11}=2\pi\sum_{k\in\mathcal{L}^{\star}}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}^{1}(2\pi k+2\pi\omega)\mathtt{d}\rho^{1}(\omega)\\ +\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}^{2}(2\pi k+2\pi\omega)\mathtt{d}\rho^{2}(\omega)\\ +\frac{1}{2}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}^{1,2}(2\pi k+2\pi\omega)\mathtt{d}(\rho^{1,2+}-\rho^{1,2-})(\omega).$
Note that $g_{\ell}^{1}$, $g_{\ell}^{2}$ and $g_{\ell}^{1,2}$ are Schwartz functions and thus so are $\widehat{g_{\ell}^{1}}$, $\widehat{g_{\ell}^{2}}$ and $\widehat{g_{\ell}^{1,2}}$ and (since $\widehat{j_{r}}$ is $L^{1}$), so are $\widehat{j_{r}\times g_{\ell}^{1}}$, $\widehat{j_{r}\times g_{\ell}^{2}}$ and $\widehat{j_{r}\times g_{\ell}^{1,2}}$ , which justifies our interversions of sums over $x\in\mathcal{L}$ and integrals against $\rho^{1}$, $\rho^{2}$, $\rho^{1,2+}$ and $\rho^{1,2-}$ as well as the use of Poisson’s formula. Defining the function $g_{\ell}:t\mapsto\chi_{\ell}(t)|t|^{2}$ and noting that it dominates $g_{\ell}^{1}$, $g_{\ell}^{2}$ and $g_{\ell}^{1,2}$ we obtain the bound
$|B_{11}|\leq 2\pi\sum_{k\in\mathcal{L}^{\star}}\int_{\Omega}j_{r}(2\pi k+2\pi\omega)g_{\ell}(2\pi k+2\pi\omega)\mathtt{d}\rho(\omega),$ (3.18)
where $\rho$ is the finite measure $\rho^{1}+\rho^{2}+\frac{1}{2}\rho^{1,2+}+\frac{1}{2}\rho^{1,2-}$.
We now distinguish between the lattice points $k=0$ and $k\in\mathcal{L}^{\star}\setminus\{0\}$.
For $k\neq 0$
@@
$B_{12}$ $=$ $\sum_{x\in\mathcal{L}}\int_{t\in\mathbb{R}^{2}}e^{-it\cdot x}j_{r}(t)\chi_{\ell}(t)\left(G^{1}(0)t_{1}^{2}+G^{2}(0)t_{2}^{2}+\frac{1}{2}(G^{1,2+}(0)-G^{1,2-}(0))t_{1}t_{2}\right)\mathtt{d}t$
$=$ $\sum_{x\in\mathcal{L}}G^{1}(0)\widehat{j_{r}\times g^{1}_{l}}(x)+G^{2}(0)\widehat{j_{r}\times g^{2}_{l}}(x)+\frac{1}{2}(G^{1,2+}(0)-G^{1,2+}(0))\widehat{j_{r}\times g^{1,2}_{l}}(x)$
$=$ $2\pi\sum_{k\in\mathcal{L}^{\star}}G^{1}(0)j_{r}(2\pi k)g_{l}^{1}(2\pi k)+G^{2}(0)j_{r}(2\pi k)g_{l}^{2}(2\pi k)$
$+\frac{1}{4}(G^{1,2+}(0)-G^{1,2+}(0))j_{r}(2\pi k)g_{l}^{1,2}(2\pi k)$
$=0.$
Indeed $g^{1}_{l}(0)=g^{2}_{l}(0)=g^{1,2}_{l}(0)=0$, and since $\chi_{\ell}$ is supported on $\mathrm{B}_{\ell}$ the other summands vanish for $\ell$ small enough (depending on $\mathcal{L}$) thus for $r$ large enough. ∎
Conclusion